OpenStudy (anonymous):

A church steeple that needs to be reshingled has the shape of a regular square pyramid. If base edges of the steeple measure 16 feet while lateral edges measure 17 feet, what is the lateral area (area to reshingle) of the steeple?

OpenStudy (cwrw238):

each face is an isosceles triangle :- 

OpenStudy (cwrw238):

the area of each triangle = 1/2 * base * perpendicular height = 1/2 * 16 * perpendicular height the height is calculated by using the pythagoras theorem h^2 = 17^2 - 16^2 multiplying this area by 4 will give you your answer

OpenStudy (anonymous):

would the answer be 4 to the square root of 33

OpenStudy (solomonzelman):

▲ . . 17 . . 17 . . . _______ . 16 If you split the bottom side, you will get 2 triangles that look like I. I . ? I . 17 I . ----- 8 but you don't know the 3rd side though. 17^2-8^2=225 and √225 = ±15 (but we are dealing with length of the side, and the length/do stance can not be negative, so just positive 15. Now.... knowing that the height is 15, and the base (of the original triangle) is 16 can you find the area ?

OpenStudy (anonymous):

would it be base times height to find the area 15x16

OpenStudy (solomonzelman):

yes, but by a triangle it is base times height -------------- 2

OpenStudy (anonymous):

duh yeah ok so its 120

OpenStudy (solomonzelman):

15 x 16 ÷ 2 15 x 8 120 I wrote it out, b/c I don't have a calculator -;( you are correct.

OpenStudy (anonymous):

lol thank you

OpenStudy (solomonzelman):

You welcome !